OpenStudy (anonymous):

The decimal number 0.75 means

OpenStudy (anonymous):

75/100 =15 / 20 = 3/ 4

OpenStudy (petiteme):

3/4

OpenStudy (anonymous):

is it there are seventy-five parts of a whole unit that has 100 parts

OpenStudy (petiteme):

yeah

OpenStudy (anonymous):

thank you i was just doubel cheking

OpenStudy (petiteme):

No prob :) Anytime :)